If 0.2F of electricity is passed through a solution of dilute sulphuric acid during electrolysis calculate the mass and volume of hydrogen gas.
1
Expert's answer
2020-06-10T07:35:19-0400

Solution:

Electrolysis of dilute sulphuric acid:

At cathode: 2H+ + 2e− ⟶ H2

​At anode: 4OH⟶ 2H2O + 2O2 + 4e−

Overall reaction: 2H2O ⟶ 2H2 + O2


​It takes a transfer of 2 electrons to form each hydrogen molecule (H2) from 2 hydrogen ions (H+).


Faraday's second law of electrolysis states that, when the same quantity of electricity is passed through several electrolytes, the mass of the substances deposited are proportional to their respective chemical equivalent or equivalent weight.


By the 2nd Law of electrolysis when 0.2 F of electricity was passed in the solution 0.2 equivalent of Hydrogen will be evolved.


Volume of 1 equivalent of Hydrogen at STP = 11.2 L

Volume of H2 = 0.2 × Volume of 1 equivalent of Hydrogen at STP

Volume of H2 = 0.2 × 11.2 = 2.24 L


Mass of H2 = 0.2 × Equivalent weight of H2

Mass of H2 = 0.2 × 1.008 = 0.2016 g = 0.2 g


Answer:

Volume of H2 is 2.24 L

Mass of H2 is 0.2 g